OpenStudy (anonymous):

simplify x^2-3x -10/x+2

OpenStudy (lgbasallote):

lol why did you repost...you could have edited the question you know :P

OpenStudy (anonymous):

oh oops :) I'll remember that next time

OpenStudy (lgbasallote):

anyway i assume this is \[\large \frac{x^2 - 3x - 10}{x+2}\] there are two possible ways this can be solved 1) simplify the numerator then cancel 2) use long division which would you like me to use?

OpenStudy (anonymous):

which ever is easier for you :)

OpenStudy (lgbasallote):

well number two is unappetizing with that \(\mathbf{long}\) word so i think i'll use number 1...okay do you know how to simplify the numerator?

OpenStudy (lgbasallote):

by simplify i mean factor out

OpenStudy (anonymous):

I horrible at math so no......can you please take me through the steps?

OpenStudy (lgbasallote):

okay....give me all the possible factors of 10

OpenStudy (anonymous):

ummm 5, 2, 1, 10

OpenStudy (lgbasallote):

good..5 and 2 then 10 and 1 can you make a -3 out of 10 and 1 using only addition/subtraction? i.e. 10 + 1 or 10 -1, etc.

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

how o.O

OpenStudy (anonymous):

you subtract 13 from 10

OpenStudy (lgbasallote):

you only use 10 and 1 :P

OpenStudy (anonymous):

oh oops didn't understand that XD no you can't

OpenStudy (lgbasallote):

it's not possible using 10 and 1... 10 + 1 = 11 10 - 1 = 9 -1 - 10 = -11 1 - 10 = -9

OpenStudy (lgbasallote):

what about 5 and 2?

OpenStudy (anonymous):

nope

OpenStudy (lgbasallote):

no? what's 2- 5?

OpenStudy (anonymous):

oh whoops XD yes then

OpenStudy (lgbasallote):

so positive 2 and negative 5 make -3 therefore (x+2)(x-5) is the factor...got it?

OpenStudy (anonymous):

ok :)

OpenStudy (lgbasallote):

another way to factor...break down the middle term... \[\large \begin{align} x^2 - 3x - 10\\ x^2 - 5x + 2x - 10\\ x(x - 5) + 2(x-5)\\ (x-5)(x+2) \end{align}\]

OpenStudy (anonymous):

ok :)

OpenStudy (lgbasallote):

anyway...rewrite your original problem \[\large \frac{(x-5)(x+2)}{x+2}\]

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

no

OpenStudy (anonymous):

oh ok well I have to go to dinner soon so can you tell me the answer?

OpenStudy (lgbasallote):

look at what i wrote a while ago...simplify the numerator then cancel...we already simplified the numerator so what's next?

OpenStudy (anonymous):

x^2-10/x+2?

OpenStudy (lgbasallote):

what? we rewrote the problem to \[\frac{(x+2)(x-5)}{x+2}\] cancel from there

OpenStudy (anonymous):

how?

OpenStudy (lgbasallote):

uhmm well look at what's common in the numerator and denominator then cancel them

OpenStudy (anonymous):

oh! so the answer is x-5!

OpenStudy (lgbasallote):

there we go

OpenStudy (anonymous):

what would x not equal though?

OpenStudy (lgbasallote):

ahh yes..that...well it's the number that will make the denominator 0..so what will make x+2 zero?

OpenStudy (lgbasallote):

hint: \[x+2 \ne 0\] find x

OpenStudy (anonymous):

-2+2=0 ....so x doesn't equal -2

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

yay! Thank you!! umm while I have you here can you answer this question: simplify x+7/x^2+4x-21 I won't be on btu can you please answer it?

OpenStudy (lgbasallote):

well i gtg :P haha and also if you have new questions post them in another post ...we're not allowed to answer two different questions in one post